# Graph is a Natural Regularization: Revisiting Vector Quantization for Graph Representation Learning

### v1

@@ Introduction
Graph VQ 94.47 4.14 60.09 1.00 1.00 10.18 13.29 10.84 20.78
EMA 91.68 5.12 55.15 1.00 1.00 11.27 9.12 6.20 14.15
AP 75.32 126.55 9.03 54.95 59.33 83.55 73.82 118.46 66.57
Reset 65.79 102.78 85.19 10.73 17.18 134.44 130.83 150.51 141.98
PT 60.57 6.17 138.98 3.78 2.94 3.10 37.65 14.49 58.99
SimVQ 40.09 23.96 38.11 37.29 40.47 45.90 16.08 42.22 21.71
HQA-GAE 130.06 164.77 93.67 166.32 114.08 98.73 92.17 89.05 72.86
RGVQ 211.69 319.09 188.17 446.02 413.10 228.82 200.93 374.51 250.79
Table 1: Mean codebook utilization in perplexity on homophilous and heterophilous graphs with codebook size $K=512$. Bold highlights the best performance. The model performance with standard deviation can be found in Appendix E.
@@
GCN 75.65 75.61 75.28 73.79 82.22 64.84 71.32 73.76
GAT 76.24 74.86 76.28 80.16 88.93 65.54 70.12 75.44
DGI 72.10 73.13 75.32 75.75 81.34 59.62 63.31 71.22
BGRL 71.20 75.29 76.53 75.44 80.66 63.95 67.09 72.67
GraphMAE 73.10 74.32 72.61 78.99 85.30 61.04 63.30 73.07
GFT 78.35 73.39 79.13 90.87 89.89 72.16 72.74 79.50
GFT + EMA 79.44 75.82 79.46 90.58 89.75 72.39 73.04 80.06
GFT + AP 79.69 74.03 78.05 89.56 89.05 71.86 71.48 79.10
GFT + Reset 80.07 73.56 79.73 91.18 88.09 72.79 71.95 79.62
GFT + PT 78.57 75.09 79.39 88.63 88.45 71.01 73.73 79.26
GFT + SimVQ 77.61 76.59 76.68 82.72 82.03 66.57 69.90 76.01
GFT + RGVQ 80.85 77.46 80.10 91.32 90.45 74.10 75.68 81.42
Table 2: Cross-domain and cross-task performance measured in Accuracy in the pre-training and fine-tuning setting. Bold highlight the best performance. Performance with standard deviation can be found in Appendix E.
Connection to Other Works. Our proposed method leverages both feature and structural information as a regularization to promote codebook utilization in graph VQ models. It serves as a general solution for codebook collapse of Graph VQ, and can be utilized as a plugin for graph models that utilize VQ, e.g., GFT (Wang et al. 2024), GQT (Wang et al. 2025), and GraphVQ (Yang et al. 2023). Moreover, it is complementary with existing collapse mitigation strategies in other domains, e.g., EMA and codebook reset. Finally, this method can be seamlessly integrated into hierarchical VQ models such as Residual VQ (Luo et al. 2024), by incorporating the regularization at each layer.
Experiments
To evaluate the generality and effectiveness of RGVQ, in this section, we conduct extensive experiments based on three key functions of Graph VQ: (1) codebook utilization, (2) transferability, and (3) serialization. First, we evaluate the codebook utilization of RGVQ on both homophilous and heterophilous graphs, aiming to verify its ability to mitigate codebook collapse. Second, we investigate the transferability. We integrate RGVQ into GFT (Wang et al. 2024), a graph foundation model that utilizes the learned codebook as pretrained graph tokens, and evaluate the performance on cross-task and cross-domain graphs. Third, we assess the serialization capability of RGVQ by examining its compatibility with sequence-based models. We use GQT (Wang et al. 2025), a transformer that takes VQ tokens as input sequences, and evaluate the performance on node classification. Detailed dataset statistics, baselines, and implementation details are provided in Appendix C. The contrastive sets in RGVQ are pre-computed before training, we provide the complexity analysis in Appendix D.
@@
GCN 75.65 75.61 73.14 96.10 93.99 75.28 53.80 91.27 79.02
GAT 76.24 74.86 72.22 96.60 94.09 76.28 55.54 90.63 77.95
GraphGPS 82.84 79.94 72.73 95.06 91.19 78.66 53.10 82.00 71.73
SGFormer 84.50 80.30 72.60 95.10 91.99 73.46 48.01 79.10 72.15
Exphomer 82.77 79.46 71.63 95.35 91.47 78.54 53.51 89.03 -
GQT 86.44 81.60 73.14 94.46 92.13 80.03 54.04 89.85 76.52
GQT + RGVQ 88.34 86.54 81.25 97.66 95.67 83.58 55.16 90.98 78.26
Table 3: Mean performance on node classification tasks. Metrics are reported in terms of ROC-AUC for Questions, and Accuracy for all other datasets. Bold indicates the best performance. Results with standard deviations are provided in Appendix E.
(a) Codebook size vs. Perplexity.
(b) Temperature $(\tau)$ vs. Perplexity.
(c) Top-K vs. Perplexity.
@@
Codebook Utilization. We evaluate codebook utilization by comparing vanilla Graph VQ and its variants: EMA (Łańcucki et al. 2020), affine parameters (AP) (Huh et al. 2023), codebook reset (Reset) (Zeghidour et al. 2021), and pretrained encoders (PT) (Zhao et al. 2024a), as well as existing SOTA VQ models: SimVQ (Zhu et al. 2024) and HQA-GAE (Zeng et al. 2025). All methods use orthogonal normalization, cosine similarity, and K-Means initialization, and are trained on the feature and link reconstruction tasks. We fix the codebook size at 512 and report perplexity in Table 1. Across every dataset, RGVQ outperforms all baselines by a clear margin and remains robust, showing that Gumbel-Softmax reparameterization combined with structure-aware regularization leads to more balanced codebook use, prevents collapse, and enables more expressive graph representations. By contrast, vanilla Graph VQ and its variants suffer from severe codebook collapse, with perplexity values as low as 1.00 in several datasets. More advanced mitigation strategies like SimVQ and HQA-GAE offer only small improvements, indicating their limited ability to fundamentally address the underlying causes of collapse in Graph VQ.
Transferability. To evaluate the effectiveness of RGVQ in learning transferrable graph tokens, we integrate it into a graph foundation model, i.e., GFT, and compare it with vanilla Graph VQ and its variants with different mitigation strategies. Moreover, we include supervised GNNs, i.e., GCN and GAT, and graph self-supervised methods, i.e., DGI (Veličković et al. 2018), BGRL (Thakoor et al. 2021), and GraphMAE (Hou et al. 2022). The supervised GNNs are trained directly on each target dataset, while the self-supervised methods and all GFT variants are pretrained on the full set of datasets and then fine-tuned per target.
@@
Table 2 reports the model performance across cross-domain and cross-task datasets. RGVQ consistently achieves the highest average performance across all tasks and datasets, surpassing both supervised and self-supervised graph models, including GFT and its variants equipped with codebook collapse mitigation strategies. These consistent gains indicate that alleviating codebook collapse is essential for improving the expressiveness and transferability of graph tokens. By encouraging more balanced and diverse codebook utilization, RGVQ learns discrete representations that generalize better across tasks and domains, positioning it as a general and robust solution for GFMs.
@@
Serialization. To further evaluate the effectiveness of RGVQ in serialization, we integrate it into the Graph Quantized Transformer (GQT), where discrete tokens serve as the input sequence to a vanilla Transformer backbone. We follow the original sequence reconstruction method (Wang et al. 2025) and compare the performance of RGVQ-enhanced GQT against the original GQT, supervised GNNs, and graph transformers, including GraphGPS (Rampášek et al. 2022), SGFormer (Wu et al. 2023), and Exphomer (Shirzad et al. 2023). Table 3 summarizes the node classification results across various benchmarks. Compared to GQT, incorporating RGVQ consistently improves classification accuracy on most datasets. These results highlight the importance of mitigating codebook collapse when learning discrete tokens for Transformer-based architectures.
Ablation Studies
@@
We further conduct ablation experiments on the contributions of the Gumbel–Softmax reparameterization and the structure-aware contrastive regularization.
Influence of the Codebook Size. We evaluate how codebook size affects the codebook utilization of RGVQ. The results are shown in Figure 4(a). Across all datasets, RGVQ consistently maintains high normalized perplexity as the codebook size increases, showing strong robustness to the choice of $K$. Notably, even with a large codebook size of 512, the model utilizes over 50% of the codebook capacity. This suggests that RGVQ can flexibly adapt to different representational granularities, and its token assignment remains stable even under large vocabulary settings.
@@
Influence of the Temperature. We investigate how the Gumbel–Softmax temperature $\tau$ affects normalized perplexity. As shown in Figure 4(b), lower temperatures, which produce distributions closer to one-hot, consistently improve codebook utilization. This indicates that, unlike some prior work (Zeng et al. 2025) that rely on temperature annealing, a relatively low and fixed temperature is sufficient to address the non-differentiability of deterministic VQ.
@@
Influence of Contrastive Samples. We examine how the number of contrastive samples in structure-aware regularization affects normalized perplexity. As shown in Figure 4(c), even a small number of contrastive samples (e.g., $K=10$ or $50$) achieves relatively high normalized perplexity, indicating effective codebook use. Increasing $K$ does not consistently lead to better utilization and may even degrade in some datasets, potentially due to added training noise.
Conclusion
@@
A. Empirical Study
Implementation Details. We provide the hyperparameters and experimental setup used in the empirical study of codebook perplexity. We jointly train the single-head VQ model and the GAT encoder using the link prediction and feature reconstruction tasks, along with the commitment loss and vocabulary loss. The task weights are set to 0.01, 100, 0.1, and 0.9, respectively. We train the model for 100 epochs and report the highest perplexity during the training process for each method and a specific codebook size $K$. For all methods, we utilize the kmeans initialization and orthogonal regulation for the codebook, with a regularization weight of 0.1. The GNN consists of 4 layers with a hidden dimension of 256. AdamW is utilized as the optimizer with a learning rate of 1e-4 and a weight decay of 1e-5. For affine parameters, we use Euclidean distance and set the codebook decay to 0.9. For codebook reset, the threshold of deadcode is set to 10. For pretraining encoder, we pretrain the GNN encoder for 50 epochs before the joint training.
Additional Experiments. In Figure 5, We provide the additional study of codebook perplexity across six datasets under varying codebook sizes, including both homophilous and heterophilous graphs. We report the study on Cora, Citeseer, and Amazon-computer in the main paper, and include PubMed, WikiCS, Amazon-Ratings, Roman-Empire, Questions, and Amazon-Photo in this appendix. Consistent with our main findings, all baseline methods exhibit severe collapse, failing to scale perplexity proportionally with the codebook capacity.
B. Proof of Co-assignment Bound
@@
Given two nodes $v_{1}$ and $v_{2}$ and their computation tree $\mathcal{T}_{v_{1}}$ and $\mathcal{T}_{v_{2}}$, let $h_{v_{1}}=\phi(\mathcal{T}_{v_{1}})$, $h_{v_{2}}=\phi(\mathcal{T}_{v_{2}})$ be their embeddings under a GNN encoder $\phi$ with the distance between node embeddings $\Delta^{L}_{v_{1},v_{2}}=\|h_{v_{1}}-h_{v_{2}}\|_{2}$. The probability that two nodes are quantized to the same codeword $p:=\mathbb{P}[z_{v_{1}}=z_{v_{2}}]$ satisfies the following lower bound:
@@
Proof.
Since the GNN encoder is based on message passing, each node’s embedding depends on information aggregated from its local neighborhood. This process can be formalized as a computation tree $\mathcal{T}_{v}^{L}$ of depth $L$, rooted at node $v$, where messages are propagated from leaf nodes up to the root. The embedding $h_{v}$ can thus be viewed as a function of this tree, i.e., $h_{v}=\phi(\mathcal{T}_{v}^{L})$. Let $x_{v}$ denote the node feature of $v$ and $\mathcal{N}_{v}$ be the set of its direct neighboring nodes, which correspond to the children of the computation tree $\mathcal{T}^{L}_{v}$. We begin by computing the embedding distance between two $L$-layer computation trees (Wang et al. 2024), whose embeddings are generated by the GNN $\phi$ with parameters $\mathbf{W}=(\mathbf{W}_{1},\mathbf{W}_{2})$, where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the transformation of the root node and its neighboring nodes, respectively. Note that both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are Lipschitz continuous and are bounded by $\|\mathbf{W}_{1}\|<\mathcal{B}_{\mathbf{W}_{1}}$ and $\|\mathbf{W}_{2}\|<\mathcal{B}_{\mathbf{W}_{2}}$. For simplicity, we assume that the GNN parameters are shared across all layers, i.e., $\mathbf{W}$ remains the same at each layer, which does not affect the validity of the proof. Then the node embedding is computed as: $h_{v}=$
$\phi(\mathcal{T}^{L}_{v})=\sigma\left(\mathbf{W}_{1}x_{v}+\mathbf{W}_{2}\rho\left(\sum_{j\in\mathcal{N}(v)}g\left(\mathcal{T}_{j}^{L-1}(\mathbf{W})\right)\right)\right),$ (16)
@@
Then, we have the lower bound of the distance between node embeddings $\Delta^{L}_{v_{1},v_{2}}=$:
$\left\|\mathcal{T}_{v_{1}}^{L}(\mathbf{W})-\mathcal{T}_{v_{2}}^{L}(\mathbf{W})\right\|_{2}$ $=\left\|\sigma\left(\mathbf{W}_{1}x_{v_{1}}+\mathbf{W}_{2}R(\mathcal{T}_{v_{1}})\right)\right$.
$\quad\left.-\sigma\left(\mathbf{W}_{1}x_{v_{2}}+\mathbf{W}_{2}R(\mathcal{T}_{v_{2}})\right)\right\|_{2}$
$\leq C_{\sigma}\left\|\mathbf{W}_{1}(x_{v_{1}}-x_{v_{2}})\right$.
$\quad\left.+\mathbf{W}_{2}\left(R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\right)\right\|_{2}$
$\leq C_{\sigma}\left\|\mathbf{W}_{1}(x_{v_{1}}-x_{v_{2}})\right\|_{2}$
$\quad+C_{\sigma}\left\|\mathbf{W}_{2}\left(R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\right)\right\|_{2}$
$\leq C_{\sigma}\mathcal{B}_{\mathbf{W}_{1}}\left\|x_{v_{1}}-x_{v_{2}}\right\|_{2}$
$\quad+C_{\sigma}\mathcal{B}_{\mathbf{W}_{2}}\left\|R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\right\|_{2}$ (17)
@@
where $\Delta^{L-1}_{v_{1},v_{2},j}=\|\mathcal{T}^{L-1}_{v_{1}}(\mathbf{W})-\mathcal{T}^{L-1}_{v_{2}}(\mathbf{W})\|_{2}$ is the distance of the embeddings between the $j$-th child of nodes $v_{1}$ and $v_{2}$. Here we assume that $v_{1}$ and $v_{2}$ have the same local structures by padding virtual nodes in their actual computation trees, as shown in Figure 6, such that such that the $j$-th child of both $v_{1}$ and $v_{2}$ always exists. These virtual nodes do not have features and do not affect the actual computation and the generality of the proof. Let $d_{l}$ denote the number of branches, i.e., number of children at the $l$-th layer, the bound in Equation 4 can be simplified as:
@@
We use $p$ to simply represent $\mathbb{P}[\|h_{v_{1}}-h_{v_{2}}\|$. According to Equation 20 and 21, we substitute $\Delta^{L}_{v_{1},v_{2}}$ with their upper bound and have:
@@
Note that this lower bound is a conservative estimate of the true co-assignment probability. This is because the bound only accounts for embedding pairs whose distance is within $\delta_{c}$, which is a sufficient condition for co-assignment. The actual probability may be higher. There may exist additional cases where $\|h_{v_{1}}-h_{v_{2}}\|>\delta_{c}$ but both embeddings still fall within the Voronoi region of the same codeword, such that $z_{v_{1}}=z_{v_{2}}$ still holds. ∎
@@
Table 4: Dataset statistics for selected datasets.
C. Experiments

### v2

@@ 6 Experiments
Graph VQ
[…]
$4±\pm1$.03
[…]
EMA
[…]
AP
[…]
$5±\pm12.6$4
[…]
Reset
[…]
PT
[…]
SimVQ
[…]
HQA-GAE
[…]
RGVQ
[…]
Table 2: Cross-domain and cross-task performance measured in Accuracy in the pre-training and fine-tuning setting. Bold highlight the best performance.
@@
GCN
[…]
74.10
GAT
[…]
76.01
GIN
[…]
69.95
DGI
[…]
71.51
BGRL
[…]
72.88
GraphMAE
[…]
72.66
GIANT
[…]
74.68
GFT
[…]
79.50
GFT + RGVQ
[…]
81.42
@@
GCN
[…]
GAT
[…]
GraphGPS
[…]
SGFormer
[…]
Exphomer
[…]
NodeFormer
[…]
GQT
[…]
GQT + RGVQ
[…]
(a) Temperature $(\tau)$.
(b) Contrastive Samples.
(c) Codebook size.
(d) Codebook diversity.
@@ 6.1 Performance Evaluation
Codebook Utilization. We evaluate codebook utilization by comparing vanilla Graph VQ and its variants: EMA (Łańcucki et al., 2020), affine parameters (AP) (Huh et al., 2023), codebook reset (Reset) (Zeghidour et al., 2021), and pretrained encoders (PT) (Zhao et al., 2024), as well as existing SOTA VQ models: SimVQ (Zhu et al., 2024) and HQA-GAE (Zeng et al., 2025). All methods use orthogonal normalization, cosine similarity, and K-Means initialization (Im & Chan, 2023), and are trained on the feature and link reconstruction tasks. We fix the codebook size at 512 and report perplexity in Table 1. Across all datasets, RGVQ outperforms all baselines by a clear margin and remains robust, showing that Gumbel-Softmax reparameterization combined with structure-aware regularization leads to more balanced codebook use, prevents collapse, and enables more expressive graph representations. By contrast, vanilla Graph VQ and its variants suffer from severe codebook collapse, with perplexity values as low as 1.00 in several datasets. More advanced mitigation strategies like SimVQ and HQA-GAE offer only small improvements, indicating their limited ability to fundamentally address the underlying causes of collapse in Graph VQ.
@@
Transferability. To evaluate the effectiveness of RGVQ in learning transferrable graph tokens, we integrate it into a graph foundation model, i.e., GFT, and compare it with vanilla Graph VQ and its variants with different mitigation strategies. Moreover, we include supervised GNNs, i.e., GCN (Zhang et al., 2022), GAT (Veličković et al., 2017), and GIN (Xu et al., 2018), and graph self-supervised methods, i.e., DGI (Veličković et al., 2018), BGRL (Thakoor et al., 2021), GraphMAE (Hou et al., 2022), and GIANT (Chien et al., 2021). The supervised GNNs are trained directly on each target dataset, while the self-supervised methods and all GFT variants are pretrained on the full set of datasets and then fine-tuned per target. Table 2 reports the model performance across cross-domain and cross-task datasets. RGVQ consistently achieves the highest average performance across all tasks and datasets, surpassing both supervised and self-supervised graph models including GFT. These consistent gains indicate that alleviating codebook collapse is essential for improving the expressiveness and transferability of graph tokens. By encouraging more balanced and diverse codebook utilization, RGVQ learns discrete representations that generalize better across tasks and domains, positioning it as a general and robust solution for GFMs.
@@
Serialization. To further evaluate the effectiveness of RGVQ in serialization, we integrate it into the Graph Quantized Transformer (GQT), where discrete tokens serve as the input sequence to a vanilla Transformer backbone. We follow the original sequence reconstruction method (Wang et al., 2025) and compare the performance of RGVQ-enhanced GQT against the original GQT, supervised GNNs, and graph transformers, including GraphGPS (Rampášek et al., 2022), SGFormer (Wu et al., 2023), Exphomer (Shirzad et al., 2023), and NodeFormer(Wu et al., 2022). Table 3 summarizes the node classification results across various benchmarks. Compared to GQT, incorporating RGVQ consistently improves classification accuracy on most datasets. These results highlight the importance of mitigating codebook collapse when learning discrete tokens for Transformer-based architectures.
6.2 Ablation Studies
@@
We further conduct ablation experiments on the contributions of the Gumbel–Softmax reparameterization and the structure-aware contrastive regularization. We also include the additional ablation studies for each proposed module in Appendix C.
@@
Influence of the Temperature. We investigate how the Gumbel–Softmax temperature $\tau$ affects perplexity. As shown in Figure 4(a), lower temperatures, which produce distributions closer to one-hot, consistently improve codebook utilization. This indicates that, unlike some prior work (Zeng et al., 2025) that rely on temperature annealing, a relatively low and fixed temperature is sufficient to regularize the codebook distributions and address the non-differentiability of deterministic VQ.
@@
Influence of Contrastive Samples. We examine how the number of contrastive samples in structure-aware regularization affects perplexity. As shown in Figure 4(b), even a small number of contrastive samples (e.g., $n=10$ or $50$) achieves relatively high perplexity, indicating effective codebook utilization. Increasing $n$ does not consistently lead to better utilization and may even degrade in some datasets, potentially due to added training noise.
Influence of the Codebook Size. We evaluate how codebook size affects the codebook utilization of RGVQ, which is defined as the ratio of utilized codebook entries to the total size of the codebook. The results are shown in Figure 4(c). Across all datasets, RGVQ consistently maintains high utilization as the codebook size increases, showing strong robustness to the choice of the codebook size. Notably, even with a large codebook size of 512, the model utilizes over 50% of the codebook capacity. This suggests that RGVQ can flexibly adapt to different representational granularities, and its token assignment remains stable even under large vocabulary settings.
Influence of Codebook Diversity. To better understand the relations between quantized diversity and the downstream performance, we evaluate how perplexity affects the node classification results. The results are shown in Figure 4(d). We observe a consistent positive correlation between perplexity and accuracy across datasets. This suggests that a more diverse codebook can capture finer-grained structural patterns, enabling the model to learn more discriminative embeddings.
7 Conclusion
@@
Appendix A Proof
@@
Given two nodes $v_{1}$ and $v_{2}$ and their corresponding computation trees $\mathcal{T}_{v_{1}}$ and $\mathcal{T}_{v_{2}}$ under a message-passing GNN encoder $\phi$, let $h_{v_{1}}=\phi(\mathcal{T}_{v_{1}})$, $h_{v_{2}}=\phi(\mathcal{T}_{v_{2}})$ be their embeddings with the distance between node embeddings $\Delta^{L}_{v_{1},v_{2}}=\|h_{v_{1}}-h_{v_{2}}\|_{2}$. The probability that two nodes are quantized to the same codeword $p:=\mathbb{P}[z_{v_{1}}=z_{v_{2}}]$ satisfies the following lower bound:
$p\geq\Bigg(1-\frac{1}{\delta_{c}}\Big[C_{1}\|x_{v_{1}}-x_{v_{2}}\|_{2}+\ C_{2}\sum_{j\in\mathcal{N}(v)}\Delta^{L-1}_{v_{1},v_{2},j}\Big]\Bigg)\geq\Bigg(1-\frac{2B_{x}}{\delta_{c}}\Big[C_{1}+\sum_{\ell=1}^{L}C_{2}^{\ell}D_{\ell}\Big]\Bigg),$ (17)
@@
Proof. Since the GNN encoder is based on message passing, each node’s embedding depends on information aggregated from its local neighborhood. This process can be formalized as a computation tree $\mathcal{T}_{v}^{L}$ of depth $L$, rooted at node $v$, where messages are propagated from leaf nodes up to the root. The embedding $h_{v}$ can thus be viewed as a function of this tree, i.e., $h_{v}=\phi(\mathcal{T}_{v}^{L})$. Let $x_{v}$ denote the node feature of $v$ and $\mathcal{N}_{v}$ be the set of its direct neighboring nodes, which correspond to the children of the computation tree $\mathcal{T}^{L}_{v}$. We begin by computing the embedding distance between two $L$-layer computation trees (Wang et al., 2024), whose embeddings are generated by the GNN $\phi$ with parameters $\mathbf{W}=(\mathbf{W}_{1},\mathbf{W}_{2})$, where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the transformation of the root node and its neighboring nodes, respectively. Note that both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are Lipschitz continuous and are bounded by $\|\mathbf{W}_{1}\|<\mathcal{B}_{\mathbf{W}_{1}}$ and $\|\mathbf{W}_{2}\|<\mathcal{B}_{\mathbf{W}_{2}}$. For simplicity, we assume that the GNN parameters are shared across all layers, i.e., $\mathbf{W}$ remains the same at each layer, which does not affect the validity of the proof. Then the node embedding is computed as:
$h_{v}=\phi(\mathcal{T}^{L}_{v})=\sigma\left(\mathbf{W}_{1}x_{v}+\mathbf{W}_{2}\rho\left(\sum_{j\in\mathcal{N}(v)}g\left(\mathcal{T}_{j}^{L-1}(\mathbf{W})\right)\right)\right),$ (18)
@@
Then, we have the lower bound of the distance between node embeddings:
$\Delta^{L}_{v_{1},v_{2}}$ $=\left\|\mathcal{T}_{v_{1}}^{L}(\mathbf{W})-\mathcal{T}_{v_{2}}^{L}(\mathbf{W})\right\|_{2}$
$=\left\|\sigma\left(\mathbf{W}_{1}x_{v_{1}}+\mathbf{W}_{2}R(\mathcal{T}_{v_{1}})\right)-\sigma\left(\mathbf{W}_{1}x_{v_{2}}+\mathbf{W}_{2}R(\mathcal{T}_{v_{2}})\right)\right\|_{2}$
$\leq C_{\sigma}\left\|\mathbf{W}_{1}(x_{v_{1}}-x_{v_{2}})+\mathbf{W}_{2}\big(R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\big)\right\|_{2}$
$\leq C_{\sigma}\left\|\mathbf{W}_{1}(x_{v_{1}}-x_{v_{2}})\right\|_{2}+C_{\sigma}\left\|\mathbf{W}_{2}\big(R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\big)\right\|_{2}$
$\leq C_{\sigma}\mathcal{B}_{\mathbf{W}_{1}}\left\|x_{v_{1}}-x_{v_{2}}\right\|_{2}+C_{\sigma}\mathcal{B}_{\mathbf{W}_{2}}\left\|R(\mathcal{T}_{v_{1}})-R(\mathcal{T}_{v_{2}})\right\|_{2},$ (19)
@@
where $\Delta^{L-1}_{v_{1},v_{2},j}=\|\mathcal{T}^{L-1}_{v_{1}}(\mathbf{W})-\mathcal{T}^{L-1}_{v_{2}}(\mathbf{W})\|_{2}$ is the distance of the embeddings between the $j$-th child of nodes $v_{1}$ and $v_{2}$. Here we assume that $v_{1}$ and $v_{2}$ have the same local structures by padding virtual nodes in their actual computation trees (Chuang & Jegelka, 2022), as shown in Figure 5, such that such that the $j$-th child of both $v_{1}$ and $v_{2}$ always exists. These virtual nodes do not have features and do not affect the actual computation and the generality of the proof. Let $d_{l}$ denote the number of branches, i.e., number of children at the $l$-th layer, the bound in Equation 20 can be simplified as:
@@
We use $p$ to simply represent $\mathbb{P}[\|h_{v_{1}}-h_{v_{2}}\|$. According to Equation 22 and 23, we substitute $\Delta^{L}_{v_{1},v_{2}}$ with their upper bound and have:
@@
Note that this lower bound is a conservative estimate of the true co-assignment probability. This is because the bound only accounts for embedding pairs whose distance is within $\delta_{c}$, which is a sufficient condition for co-assignment. The actual probability may be higher. There may exist additional cases where $\|h_{v_{1}}-h_{v_{2}}\|>\delta_{c}$ but both embeddings still fall within the Voronoi region of the same codeword, such that $z_{v_{1}}=z_{v_{2}}$ still holds.
Table 4: Dataset statistics for selected datasets.
@@
Appendix B Experimental Setup
B.1 Dataset